OpenStudy (anonymous):

proving identities: secx+tanx=1/secx-tanx ? how?

OpenStudy (abb0t):

HINT: \(sec(x) = \frac{1}{cos(x)}\) and \(tan(x) = \frac{sin(x)}{cos(x)}\)

OpenStudy (anonymous):

first, decide which side is more complicated and change it to as @abbot said

OpenStudy (anonymous):

well i know the identities but can't really solve it sir.

OpenStudy (abb0t):

I'm looking at it and this identity looks a bit strange. r u sure that's the identity given?

OpenStudy (abb0t):

is it \(\frac{1}{sec(x)-tan(x)}\) or \(\frac{1}{sec(x)} - tan(x)\)

OpenStudy (anonymous):

\[\frac{1}{a}+\frac{b}{a}=\frac{1}{\frac{1}{a}-\frac{b}{a}}\]

OpenStudy (anonymous):

yes sir. even I am confused if this identity really can be solved. the book says, prove the identities. there's no more instruction about it is an identity or not.

OpenStudy (anonymous):

the first one sir abbot

OpenStudy (anonymous):